Consider a producer facing the following cost function MC(q,w)=100+q Assuming the fixed cost are 5000 and rhe output price is 100 find maximum profit


1
Expert's answer
2021-02-02T10:30:46-0500

By the definition of the total revenue, we get:


"TR=PQ=100Q."


Let's find the marginal revenue:


"MR=\dfrac{dTR}{dQ}=\dfrac{d}{dQ}(100Q)=\$100."


The profit maximizing level of output will be when MR=MC:


"100+Q=100,""Q=0."

Let's write the TC function:


"TC=FC+VC=5000+\int MC(Q)dQ,""TC=5000+\int (100+Q)dQ=5000+100Q+0.5Q^2."

By the definition of the profit, we have:


"\pi=TR-TC=100Q-5000-100Q-0.5Q^2,""\pi=100\cdot0-5000-100\cdot0-0.5\cdot(0)^2=-\$5000."